OpenStudy (anonymous):

what is the 7th term of the geometric sequence 3/128, 3/16 ,3/2 ,12

OpenStudy (lochana):

you can say\[\frac{3}{2^{(7-3n)}}\] for nth term of the sequence

OpenStudy (lochana):

for example, put n=0 if you want first term

OpenStudy (lochana):

put n=1, if you want 2nd term

OpenStudy (lochana):

put n = 2 if you want 3rd term

OpenStudy (lochana):

and put n = 6 for the 7th term

OpenStudy (lochana):

get it?

OpenStudy (anonymous):

yes

OpenStudy (lochana):

\[7th \ term = \frac{3}{2^{7 - 6*3}}\]

OpenStudy (lochana):

\[\frac{3}{2^{-11}}\]\[3*2^{11}\]

OpenStudy (lochana):

3*2048 = 6144

OpenStudy (anonymous):

i got 6144

OpenStudy (lochana):

nice:) bye